### v1

# Resolving Build Conflicts via Example-Based and Rule-Based Program Transformations

Abstract.
Merge conflicts often arise when developers integrate changes from different software branches. The conflicts can result from overlapping edits in programs (i.e., textual conflicts), or cause build and test errors (i.e., build and test conflicts). They degrade software quality and hinder programmer productivity. While several tools detect build conflicts, few offer meaningful support for resolvinntional cases like those caused by method removal. To overcome limitations of existing tools, we introduce BuCoR (BUild COnflict Resolver), a new conflict resolver. BuCoR first detects conflicts by comparing three versions related to a merging scenario: base $b$, left $l$, and right $r$. To resolve conflicts, it employs two complementary strategies: example-based transformation (BuCoR-E) and rule-based transformation (BuCoR-R). BuCoR-R applies predefined rules to handle common, well-understood conflicts. BuCoR-E mines branch versions ($l$ and $r$) for exemplar edits applied to fix related build errors. From these examples, it infers and generalizes program transformation patterns to resolve more complex conflicts.
We evaluated BuCoR on 88 real-world build conflicts spanning 21 distinct conflict types. BuCoR generated at least one solution for 65 cases, and correctly resolved 43 conflicts. We observed that this hybrid approach—combining context-aware, example-based learning with structured, rule-based resolution—can effectively help resolve conflicts. Our research sheds light on future directions of more intelligent and automated merge tools.
software merge, build conflict, static analysis, example-based, rule-based, program transformation, example mining, pattern inference
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software evolution
1. Introduction
Version control systems (VCSs) like Git are popularly used in collaborative software development. With VCSs, programmers create and work on separate branches for feature addition, bug fixing, or feature improvement. Periodically, they merge branches into a primary branch, to integrate the edits applied to distinct branches into one program version. In such a scenario, merge conflicts can happen if the edits from different branches are incompatible. It is challenging and time-consuming to properly handle conflicts. Prior work shows that developers often spend hours or days detecting and resolving conflicts before correctly merging branches (Kasi and Sarma, 2013).
As illustrated in Figure 1, a typical merging scenario involves five program versions: two branch versions $l$ and $r$ whose edits need to be merged, base version $b$—the common origin of both branches, the automatically merged version $A_{m}$ produced by git-merge when it naïvely integrates branch edits textually, and the manually merged version $m$ that developers create based on $A_{m}$. Among the various possible conflicts between $l$ and $r$, build conflicts refer to the incompatible edits whose naïve integration triggers build errors in the resulting merged version. As shown in Figure 3, because $l$ adds a call to $m()$ while $r$ renames that method, the co-application of both edits can cause a compilation error of unresolved method reference. Namely, the newly added method call $m()$ is not associated with any defined method.
Figure 1. A typical merging scenario can involve up to five program versions
Although tools were created to detect build conflicts (Towqir et al., 2023; Wuensche et al., 2020a; Shen et al., 2019), there is rare automatic support to resolve such conflicts. In particular, Gmerge (Zhang et al., 2022) relies on Clang compiler message to locate conflict-related changes in branches, and applies k-shot learning with GPT-3 to suggest symbol renaming for conflict resolution. MrgBldBrkFixer (Sung et al., 2020) starts with a naïvely merged version $A_{m}$, looks for build errors due to failed resolution of symbols, identifies symbol renaming changes in branches-to-merge, extracts related patches in either branch, and similarly applies those patches to fix build conflicts. These tools suffer from three limitations:
Figure 2. A build conflict due to the updated def of m() by $r$, and added use by $l$
Figure 3. BuCoR leverages Bucond (Towqir et al., 2023) to detect build conflicts, and employs two complimentary strategies to resolve conflicts
  1. (1)
    Both tools adopt compiler errors to locate conflicts; if compilers fail to locate errors, neither tool works.
  2. (2)
    They only handle conflicts related to symbol renaming (Figure 3), but fail to resolve many other conflicts, such as those caused by class-hierarchy change, method addition/deletion, and parameter addition/deletion.
  3. (3)
    They suggest simple edits of symbol renaming to adapt uses of renamed entities, but fail to suggest complex edits that systematically change uses and surrounding context.
To overcome those limitations, we introduce BuCoR (BUild COnflict Resolver), a novel resolver of build conflicts. As shown in Figure 3, given the three program versions ($l$, $b$, $r$) related to a merging scenario, BuCoR first applies Bucond (Towqir et al., 2023) to reveal conflicting edits between branches. As Bucond detects conflicts via static analysis instead of automatic build, BuCoR eliminates dependencies on compilers. BuCoR also adopts git-merge to generate a naïvely merged version $A_{m}$, extending Bucond to (1) create a graphical representation for $A_{m}$ and (2) relate that graph with those of the three input program versions. Based on the generated conflict report and four-way interconnected graph, BuCoR opportunistically applies two complementary strategies: BuCoR-E and BuCoR-R.
Example-based transformation (BuCoR-E): Prior work shows that many build conflicts are caused by mismatches between revised or removed def (short for “definition”), and newly introduced use of the same program entities (e.g., classes and methods) (Shen et al., 2022; Da Silva et al., 2022). Inspired by this insight, BuCoR-E identifies the def-change responsible for each reported conflict, locates the owner branch, and mines that branch for any exemplar edit applied to adjust corresponding uses. It then infers a program transformation pattern from that edit. For each use-addition responsible for the reported conflict, BuCoR-E tentatively establishes context matching between the pattern and code in $A_{m}$. If a full or partial match is found, BuCoR-E customizes and applies the entire or partial pattern for conflict resolution.
Rule-based transformation (BuCoR-R): Prior studies reveal patterns frequently applied to resolve certain kinds of build conflicts (Shen et al., 2022; Da Silva et al., 2022). Following those studies, we defined and implemented 16 resolution rules/patterns in BuCoR-R. Given a conflict, BuCoR-R searches its rule set for an applicable rule; when a rule is found, BuCoR-R customizes that rule for edit application.
To evaluate the tool effectiveness, we applied BuCoR to 88 real-world build conflicts in 30 open-source projects. BuCoR correctly resolved 43 conflicts. BuCoR-E and BuCoR-R separately generated resolutions for 28 and 51 conflicts; 21 and 29 of those resolutions separately match developers’ resolutions recorded in version history. All these numbers demonstrate BuCoR’s effectiveness in handling conflicts. To sum up, we made the following contributions:
  • We created BuCoR, a new static analysis-based resolver of build conflicts, to novelly combine example-based transformation with rule-based transformation. Unlike prior work, BuCoR does not adopt compilers to detect conflicts.
  • We explored BuCoR-E, an advanced example-based approach of conflict resolution. Different from existing tools, it applies program dependency analysis to derive resolution patterns from exemplar edits, establishes context matching for partial/full pattern customization as well as application, and ranks candidate resolutions to suggest the best one.
  • We explored BuCoR-R, a rule-based approach to resolve frequently occurred conflicts in 16 conventional ways. No prior work implements such an approach.
  • We systematically evaluated BuCoR with a dataset of 88 conflicts that span 21 types, and observed novel phenomena.
Table 1. A motivating example of build conflict, whose resolution requires context-specific edits more than symbol renaming
Edits from the branches-to-merge
(a) Conflicting edits between branches (b) BuCoR mines branch edits for exemplar edit E to adapt code
Changes in $\boldsymbol{l}$ (responsible def updates): In TypeSerializerConfig.java, names of the file, Java class, and constructor are all updated to SerializerConfig. Changes in $\boldsymbol{r}$ (responsible use introduction): In a newly added file XmlClientConfigBuilder.java, a method is defined to access TypeSerializerConfig + private void handleSerializers(Node node, SerializationConfig serializationConfig) { + + if (”type-serializer”.equals(name)) { + TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); + typeSerializerConfig.setClassName(value); + final String typeClassName = getAttribute(child, ”type-class”); + typeSerializerConfig.setTypeClassName(typeClassName); + serializationConfig.addTypeSerializer(typeSerializerConfig); + } } Edit example in $\boldsymbol{l}$ to adapt usage of TypeSerializerConfig: private void handleSerializers(Node node, …) { - if(”type-serializer”.equals(name)){ - TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); - typeSerializerConfig.setClassName(value); + if(”serializer”.equals(name)){ + SerializerConfig serializerConfig = new SerializerConfig(); + serializerConfig.setClassName(value); final String typeClassName=retrieveAttribute(child, ”type-class”); - typeSerializerConfig.setTypeClassName(typeClassName); - serializationConfig.addTypeSerializer(typeSerializerConfig); + serializerConfig.setTypeClassName(typeClassName); + serializationConfig.addSerializerConfig(serializerConfig);
Tool-generated edits that are applicable to $A_{m}$, to resolve conflicts
(c) A naïve resolution producible by existing tools and BuCoR-R (d) A more comprehensive resolution produced by BuCoR-E
private void handleSerializers(Node node, …) { if (”type-serializer”.equals(name)) { -   TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); +   SerializerConfig typeSerializerConfig = new SerializerConfig(); typeSerializerConfig.setClassName(value); final String typeClassName=getAttribute(child, ”type-class”); typeSerializerConfig.setTypeClassName(typeClassName); serializationConfig.addTypeSerializer(typeSerializerConfig); } } private void handleSerializers(Node node, …) { - if (”type-serializer”.equals(name2)) { -  TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); -  typeSerializerConfig.setClassName(value); + if (”serializer”.equals(name2)) { +  SerializerConfig serializerConfig = new SerializerConfig(); +  serializerConfig.setClassName(value); final String typeClassName=getAttribute(child, ”type-class”); -  typeSerializerConfig.setTypeClassName(typeClassName); -  serializationConfig.addTypeSerializer(typeSerializerConfig); +  serializerConfig.setTypeClassName(typeClassName); +  serializationConfig.addSerializerConfig(serializerConfig); } }
2. A Motivating Example
To facilitate discussion, this section introduces a running example we crafted based on a real-world project hazelcast (haz, 2022). As shown in Table 1(a), a merging scenario has $l$ rename a class and $r$ insert code to use the original class. The naïve integration of these edits can trigger a build error as the newly introduced uses refer to a nonexistent class def; thus, a build conflict occurs.
To resolve such a conflict, existing tools simply update class references by replacing TypeSerializerConfig with SerializerConfig (Table 1(c)). However, such a replacement is insufficient, as the context still has variable typeSerializerConfig and literal “type-serializer” match the original class usage. Consequently, after existing tools update class references, developers need to manually replace defs and uses of related variables/literals, to ensure consistent updates and prevent semantic conflicts. Furthermore, when such class uses are introduced at multiple places, developers have to go over all places to manually apply those edits again and again, which process is tedious and error-prone.
To overcome the limitations of existing tools and better help developers, we introduce BuCoR—a hybrid approach to combine example-based resolution with rule-based resolution. The example-based resolution BuCoR-E is derived from our insight, on the association between build errors in branches and build conflicts in the merged version. Basically, many build errors are caused by mismatches between def and use of the same program entities; many build conflicts are caused by mismatches between modified def and newly introduced use of entities (Shen et al., 2022; Da Silva et al., 2022). If on either branch, developers resolved def-use mismatches by applying specialized edits; then they are likely to reapply the same or similar edits to resolve conflicts that show the same kind of mismatches in software merge.
For each reported conflict, BuCoR-E locates the responsible def-change, and mines the contributing branch for developers’ exemplar edit $E$ applied to adjust existing uses of the changed entity. As shown in Table 1(b), the edit example is extracted from $l$ by BuCoR, because that edit adjusts usage of TypeSerializerConfig—the renamed class. The edit $E$ not only updates references to the old class/constructor, but also revises a related variable typeSerializerConfig, a literal “type-serializer”, and a method call addTypeSerializer(...). BuCoR-E then derives a transformation pattern $P$ from $E$, by abstracting away irrelevant edit detail and/or program context.
For the responsible use-introduction of each conflict, BuCoR-E tentatively establishes context matching between $P$ and the edit location. If the matching succeeds, BuCoR-E customizes $P$ by generating edit operations with respect to the matched context. BuCoR then applies those operations to transform code. As shown in Table 1(d), the context-to-handle is different from the original edit context (see Table 1(b)): it uses a different variable (name2 vs. name), and invokes a different method (getAttribute(...) vs. retrieveAttribute(...)). Albeit the differences, BuCoR still resolves the conflict by mimicking developers’ coding practices.
In addition to BuCoR-E, our tool also integrates a rule-based transformation approach BuCoR-R. This is because there are scenarios where $l$ and $r$ do not have exemplar edit $E$ in response to def-changes, making BuCoR-E less useful. BuCoR-R can opportunistically suggest resolutions when BuCoR-E does not work, to mitigate the limitation of example-based transformation. For the merging scenario described above, BuCoR suggests two alternative resolutions for developers to review (see Table 1(c)–(d)).
3. Approach
As shown in Figure 3, BuCoR has four components: conflict detection, graph construction, BuCoR-E, and BuCoR-R. This section explains each of them in detail.
3.1. Conflict Detection (Bucond (Towqir et al., 2023))
[…]
For implementation, Bucond uses JavaParser to parse source code, and adopts JGraphT (jgr, 2025) to construct and analyze PEGs.
3.2. Graph Construction
To facilitate users’ conflict comprehension and our tool’s resolution placement, we extended Bucond to construct a four-way interconnected graph. Our extension involves two parts: (1) generating a merged version $A_{m}$ and (2) comparing its PEG with those of given program versions. As a first step, BuCoR uses the widely used tool git-merge (git, 2019) to create a naïvely merged version $A_{m}$. While git-merge can detect textual conflicts—cases where multiple branches apply divergent edits to the same code fragment, it does not detect or resolve build conflicts. The $A_{m}$ it produced offers a program context for which BuCoR later suggests resolution edits.
In Step 2, to relate $A_{m}$ with $l$ and $r$, BuCoR also creates a PEG for $A_{m}$, and matches this graph against the PEGs of other versions. As shown in Figure 4, we denote the PEGs as $G_{b}$, $G_{l}$, $G_{r}$, $G_{Am}$, corresponding to the base, left, right, and naïvely merged versions. The differences between $G_{l}$ and $G_{b}$, and between $G_{r}$ and $G_{b}$, are represented as $\Delta(G_{l},G_{b})$ and $\Delta(G_{r},G_{b})$; they capture entity-level edits like entity addition/deletion/update, and relation addition/deletion. They are computed by Bucond.
We use $\cap(G_{Am},G_{l})$ and $\cap(G_{Am},G_{r})$, to separately denote intersections between $G_{Am}$ and $G_{l}$, and between $G_{Am}$ and $G_{r}$. They represent structural commonality between versions, serving as anchors to align program context as well as edits between $A_{m}$, $l$, and $r$. BuCoR recognizes such commonality using the graph-matching algorithm from Bucond. Intuitively, given two graphs under comparison, the algorithm first identifies exact node matches based on entity types and fully qualified names (FQNs). For nodes that remain unmatched, it ambiguously matches nodes based on the similarity of their internal code implementation and surrounding context. The identified common elements across program versions enable BuCoR to accurately position edits during conflict resolution.
Figure 4. Four-way interconnected graph for the motivating example in Table 1
3.3. Example-based Transformation (BuCoR-E)
To resolve a conflict, BuCoR-E operates in three phases. It first mines branches for edit example(s), infers a transformation pattern from each example, and then customizes as well as applies those patterns to revise $A_{m}$.
3.3.1. Phase I. Example Mining
Given a conflict, BuCoR-E identifies the responsible def-change, mines the branch contributing that change, and looks for any exemplar edit which adapts entity uses and surrounding code for that change. In our motivating example, because $l$ renames class TypeSerializerConfig and its constructor, it is the contributor branch of def-change. By traversing entity-level edits of $l$$\Delta(G_{l},G_{b})$, BuCoR-E locates changes of the corresponding class/constructor references. Namely, if an entity $\boldsymbol{e}$ uses TypeSerializerConfig in $b$ but uses SerializerConfig in $l$, then the entity $e$ contains an edit example. Certainly, if multiple entities adapt their usage to the same def-change, BuCoR-E considers all these entities to have relevant examples.
To extract and represent each edit example, BuCoR-E applies an off-the-shelf syntactic differencing tool GumTree (Falleri et al., 2014) to the base and branch versions of $e$, to generate an edit script of Abstract Syntax Tree (AST). The script may have four kinds of operations:
[…]
For our motivating example, the script extracted from the located exemplar edit includes eight operations: each operation updates a node to replace a string literal, type usage, constructor/method usage, or variable usage (see Figure 5).
Figure 5. From the edit example shown in Table 1(b), eight AST edit operations are extracted (see ①–⑧)
3.3.2. Phase II. Pattern Inference
In each example $E$, not every edit operation was applied to adapt uses for a given def-change, and not all unchanged code is relevant to those adaptive changes. To infer a minimal transformation pattern $P$, BuCoR-E takes two steps: edit refinement and context refinement.
Step 1 of Phase II: Edit Refinement.
BuCoR-E extracts edit operations related to uses for a given def-change, but abstracts away irrelevant operations. Specifically, if a def-change revises an existing field or method, BuCoR-E considers all accesses of that field/method as uses. If a def-change revises a class, BuCoR-E considers (1) all accesses to that class and (2) all accesses to the variables instantiated with that class as uses. For our motivating example, as shown in Figure 5, edit operation ② is relevant as it updates usage of the changed class TypeSerializerConfig; ④ is relevant as it updates usage of the changed constructor TypeSerializerConfig; operations ③, ⑤, ⑥, and ⑧ are relevant as they all update usage of variable typeSerializerConfig—an instance of TypeSerializerConfig.
Starting with the initial set of use-related operations ($E_{0}$), BuCoR-E further includes operations that occur within the same statement(s) as any operation in $E_{0}$. For instance, operation ⑦ gets incorporated because it appears in the same statement as ⑧. Our rationale is that when two edit operations are applied to the same statement, they are very likely to be semantically related. To conservatively preserve the completeness of extracted edits, we include all operations co-applied with use-related operations in the same statements. We denote the expanded operation set with $E_{1}$.
As shown in Figure 5, BuCoR-E locates edited nodes for the entire edit script in base $b$, identifies the subset of edited nodes covered by $E_{1}$, and conducts static analysis to find other edited nodes directly depended on by the subset. BuCoR-E extracts two kinds of statement-level dependencies:
  • Control Dependency: $x$ depends on $y$, if whether or not $x$ is executed depends on the execution outcome of $y$.
  • Data Dependency: $x$ depends on $y$ if $x$ uses a variable defined by $y$.
Such dependencies are essential to reveal changes that constrain any use-change. Namely, without the operations on which $E_{1}$ depends, $E_{1}$ can wrongly modify semantics and introduce errors like accessing undefined variables. BuCoR-E adopts WALA (wal, 2024), a widely used program analysis tool, to conduct dependency analysis. By extracting use-related operations and any operations they depend on, BuCoR-E ensures to abstract away non-essential co-applied edit. For our motivating example, ① also gets included, as the if-condition controls whether or not edited nodes in $E_{1}$ get executed.
Step 2 of Phase II: Context Refinement.
BuCoR-E extracts the minimal subtree that covers all refined operations, to abstract away non-essential context and derive a conflict-resolution pattern. Listing 1 shows the pattern BuCoR-E infers from Table 1(b). Compared with the original example, this pattern is more concise. It has all operations necessary to adapt the usage of an updated class and its constructor. Meanwhile, it removes surrounding program context of the if-statement, and an unchanged statement in that structure: final String typeClassName=getAttribute(child, "type-class");.
Listing 1: A pattern inferred from the mined edit example
- if(”type-serializer”.equals(name)){
- TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig();
- typeSerializerConfig.setClassName(value);
- typeSerializerConfig.setTypeClassName(typeClassName);
- serializationConfig.addTypeSerializer(typeSerializerConfig);
+ if(”serializer”.equals(name)){
+ SerializerConfig serializerConfig = new SerializerConfig();
+ serializerConfig.setClassName(value);
+ serializerConfig.setTypeClassName(typeClassName);
+ serializationConfig.addSerializerConfig(serializerConfig);
3.3.3. Phase III: Pattern Application
To resolve conflicts, BuCoR-E enumerates patterns inferred for each conflict-responsible entity use, tentatively matching each pattern $P$ with the program context to see whether the pattern can be customized and applied. This phase has two steps: context matching and AST rewriting.
Input: Pattern $P$, method-to-edit $M$, changed entity $e$
Output: Node matches between the ASTs: $W$
/* Find a match for the last usage of $e$/
0.1 Locate a statement $s_{p}\in P$, which has the last usage of $e$
0.2 Search among $M$ for a statement $s_{m}$ that best matches $s_{p}$
0.3 if match is found then
0.4  $W.add(s_{p},s_{m},matchingScore)$
  /* Use the matched nodes as anchors to guide tentative matches for siblings */
0.5  foreach sibling statement $l_{p}$ coming before $s_{p}$ do
0.6     foreach sibling statement $l_{m}$ coming before $s_{m}$ do
0.7        Search for a best match in a backward manner
0.8        if a best match is found then
0.9           $W.add(l_{p},l_{m},bestScore)$
0.10
0.11
0.12 foreach sibling statement $l_{p}$ coming after $s_{p}$ do
0.13     foreach sibling statement $l_{m}$ coming after $s_{m}$ do
0.14        Search for a best match in a forward manner
0.15        if a best match is found then
0.16           $W.add(l_{p},l_{m},bestScore)$
0.17
0.18
 /* Compare parent nodes of $s_{p}$ and $s_{m}$/
0.19  Locate the parent node of $s_{p}$, $p_{p}\in P$, if there is any
0.20  Locate the parent node of $s_{m}$, $p_{m}\in M$, if there is any
0.21  Check whether $p_{m}$ matches $p_{p}$
0.22
Repeat 1.3–1.17 by treating $p_{p}$ as $s_{p}$ and treating $p_{m}$ as $s_{m}$, until no more parent can be found or the parent match fails
Algorithm 1 The context-matching algorithm
Step 1 of Phase III: Context Matching.
Because Phase II applies backward dependency analysis to use-related edited nodes to infer $P$, the resulting contextual AST typically includes critical nodes (i.e., edited nodes using entity $e$) as leaf nodes, and noncritical ones (i.e., unchanged nodes or edited nodes not using $e$) as inner nodes. Our goal is to adapt usage of $e$. Thus, to decide whether an inferred pattern is applicable, we prioritize the match search for critical nodes. Namely, if critical nodes do not match, a pattern is inapplicable even if noncritical nodes match. If critical nodes match, a pattern is at least partially applicable even if higher-level noncritical ones do not match. We designed an algorithm to perform use-centric, bottom-up context matching.
As illustrated by Algorithm 1, given a pattern $P$, a method $M$ whose entity use causes a conflict, and the changed entity $e$, BuCoR-E tentatively matches nodes in a bottom-up manner with the best effort. Specifically, it first locates a statement $s_{p}$ in $P$ that contains the last usage of $e$, to search among $M$ for a statement $s_{m}$ best matching $s_{p}$ (see 1.1–1.2). To compare statements, we check their AST node types and values to calculate a matching score as below:
  • Type Match: Two nodes, $x$ and $y$, have a type match if their node types are identical. Such a match increments the initial matching score (i.e., 0) by 1. Type match enables patterns to get applied to the context (1) distinct from the original edit example, but (2) preserving the structural consistency.
  • Value Match: Two nodes, $x$ and $y$, have a value match if their strings have a 3-gram similarity score greater than 0.618. Here, the string of a simple statement (e.g, MethodInvocation) includes all content of that statement; the string of a complex statement (e.g., conditional or loop structure) only includes the structure header (e.g., if-condition or for-loop header). A successful value match increments the overall matching score by the calculated similarity value. This threshold-based approach allows similar nodes to match, even though they are not identical. Such a flexibility also increases the applicability of inferred patterns.
Particularly, if $x$ and $y$ are identical, they have a type match and the 3-gram similarity is 1, so the overall matching score is 2.
Once a best match of $s_{p}$ is found (with score $>$ 1.618), BuCoR-E treats the pair $(s_{p},s_{m})$ as anchor points to guide node matching for their siblings (see 1.5–1.14). Specifically for siblings preceding $s_{p}$, it enumerates nodes in reverse order: it first compares $s_{p}$’s nearest sibling against all siblings preceding $s_{m}$ for a best match; it then proceeds to match $s_{p}$’s more distant siblings against any remaining unmatched nodes, moving backward from the most recently matched one. Similarly, for siblings following $s_{p}$, BuCoR-E matches nodes in forward order.
After searching best matches for all sibling nodes of $s_{p}$, the process continues to tentatively match the parent node of $s_{p}$ with that of $s_{m}$. If this match succeeds, BuCoR-E repeats 1.3–1.17 in Algorithm 1, using the newly matched blocks as anchors for further match. The procedure continues until no more parent node in $P$ can be found or the parent match fails. In the end, the algorithm outputs best matches of nodes, and their matching scores.
If Phase II infers multiple patterns from edit examples, these patterns may be all applicable to the method-to-edit $M$. To find the best one, we defined two ways to score the overall context matching between each pattern and $M$, based on the output of Algorithm 1:
  • Sum of Matching Scores ($\Sigma_{m}$): Adding up the scores of all best matches.
  • Count of Exact Matches ($C_{m}$): Counting the number of exact matches (i.e., matching score = 2).
We believe that the greater the contextual similarity between a candidate pattern $P_{i}$ and $M$, the easier it is to apply that pattern and the more likely the resulting version is correct. Thus, among all applicable patterns, we select the one with the highest $\Sigma_{m}$. If multiple candidates have a tie, we choose the one with the highest $C_{m}$, ensuring a closer semantic match between that pattern and $M$.
In our motivating example, the method-to-edit (Table 1(c)) shares a highly similar context with the inferred pattern (Listing 1): $\Sigma_{m}=9.94$ and $C_{m}=4$. All five edited statements in the pattern sequentially find best matches within the method: four statements match exactly, and the if-condition has a highly similar counterpart.
Step 2 of Phase III: AST Rewriting.
To apply the selected pattern to $M$, BuCoR-E manipulates the AST of $M$ based on context-matching results, and pretty-prints the updated AST to suggest a resolved version to developers. Although BuCoR-E matches context using statement-level similarity calculation, it can manipulate ASTs at a finer granularity (e.g., updating a symbol). This is because the edit operations inferred from examples by GumTree are finer-grained; reapplying these operations in new context minimizes the modification and helps prevent unwanted changes. Furthermore, if the selected pattern only has partial context to match $M$, only the edit operations corresponding to the matched part get applied; the remaining ones corresponding to unmatched part are not applied. In this way, BuCoR-E resolves conflicts with the best effort, to maximize its applicability and conflict-resolution capability.
3.4. Rule-based Transformation (BuCoR-R)
BuCoR-R resolves conflicts via rule-based program transformations. To define practical resolution rules or patterns, we manually inspected the rules and data observed by prior empirical studies (Shen et al., 2019; Da Silva et al., 2022), and expanded those rule sets to generalize rules across similar but different conflict types. Table 2 shows the resulting 16 conventional patterns we implemented in BuCoR-R; each pattern resolves one type of frequent conflicts. Among the 16 conflict types, only 6 types involve symbol renaming (C1, C6, C7, C11, C13, C15). As prior work only resolves conflicts caused by symbol renaming, they cannot handle 10 of the conflict types BuCoR-R focuses on.
Table 2. The 16 resolution rules used in BuCoR-R
Idx Conflict Type Resolution Pattern
C1 Class: rename def vs. add use Update the added use
C2 Class: add method def in super vs. add sub class Update method def in sub to match super
C3 Class: change a method’s parameter list in super vs. add sub class Update method def in sub to match super
C4 Class: change a method’s return type in super vs. add sub class Update method def in sub to match super
C5 Import: remove def vs. add use Re-add the def (i.e., add back the removed entity import)
C6 Package: rename def vs. add use Update the added use (i.e., update import with the new package name)
C7 Interface: rename def vs. add use Update the added use
C8 Interface: add method def in super vs. add class to implement the super Add/update method def in class to override the new method in super
C9 Interface: change a method’s parameter list in super vs. add class to implement the super Update parameter list in class to match super
C10 Interface: remove method def in super vs. add class to implement the super Remove method def in class to match super
C11 Interface: rename a method def in super vs. add class to implement the super Rename method def in class to match super
C12 Interface: change a class to implement the interface vs. change a method’s return type in the class Update return type of method in class to match super
C13 Field: rename def vs. add use Update the added use
C14 Field: add def vs. add def Remove redundant field definition
C15 Method: rename def vs. add use Update the added use
C16 Method: add def vs. add def Remove redundant method definition
Table 3. 21 conflict types of the 88 conflicts in our dataset
Idx Conflict Type # of Conflicts
C1 Class: rename def vs. add use 7
C2 Class: add method def in super vs. add sub class 2
C3 Class: change a method’s parameter list in super vs. add sub class 2
C4 Class: change a method’s return type in super vs. add sub class 1
C5 Import: remove def vs. add use 5
C6 Package: rename def vs. add use 2
C8 Interface: add method def in super vs. add class to implement the super 6
C9 Interface: change a method’s parameter list in super vs. add class to implement the super 1
C10 Interface: remove method def in super vs. add class to implement the super 1
C11 Interface: rename a method in super vs. add class to implement the super 2
C12 Interface: change a class to implement the super vs. change a method’s return type in the class 9
C14 Field: add def vs. add def 3
C15 Method: rename def vs. add use 8
C16 Method: add def vs. add def 2
C17 Class: remove def vs. add use 10
C18 Constructor: change the parameter list vs. add use 5
C19 Field: change a field’s type vs. add use 1
C20 Field: remove def vs. add use 8
C21 Method: change the parameter list vs. add use 4
C22 Method: change the return type vs. add use 1
C23 Method: remove def vs. add use 8
BuCoR-R has a Rule Applier to opportunistically resolve conflicts based on predefined rules. Namely, for each reported conflict, BuCoR-R tentatively matches the conflict against documented conflict types. If a match is found, Rule Applier applies helper functions to the responsible entity’s def or use, to rewrite the AST nodes.
Table 4. The experiment result of BuCoR
# of Resolutions Generated # of Correct Resolutions Coverage (C) Accuracy (A)
BuCoR-E 28 21 32% (28/88) 75% (21/28)
BuCoR-R 51 29 58% (51/88) 57% (29/51)
BuCoR 79 50 74% (65/88) 63% (50/79)
4. Experiment
To evaluate BuCoR, we defined three research questions (RQs):
  • RQ1: How often can BuCoR generate resolutions?
  • RQ2: What percentage of generated solutions are correct?
  • RQ3: What is the effectiveness comparison between BuCoR-E and BuCoR-R?
We conducted the experiment on a Windows machine with AMD Ryzen 9 8945HS @4.00GHz and 16 GB memory. We did not empirically compare BuCoR with Gmerge (Zhang et al., 2022) or MrgBldBrkFixer (Sung et al., 2020), as they are close-sourced tools targeting C/C++ programs while BuCoR focuses on Java code. The following subsections describe our dataset, evaluation metrics, and results.
4.1. Dataset
We constructed our evaluation dataset by reusing the datasets of prior work (Towqir et al., 2023; Da Silva et al., 2022; Shen et al., 2019) with our best effort. Given a merging scenario with at least one known build conflict, we decided whether it is reusable based on the following criteria:
  • a)
    Both $l$ and $r$ build successfully.
  • b)
    The automatically merged version $A_{m}$ output by git-merge does not contain any textual conflict.
  • c)
    The build conflict is detectable by Bucond—BuCoR’s conflict detection module.
We ended up with a dataset of 88 build conflicts from 55 merging scenarios, which were mined from in total 30 open-source Java repositories. Most of the cases we filtered out do not satisfy a) or b). As shown in Table 3, the 88 conflicts belong to 21 types: 14 of the types overlap with those mentioned in BuCoR-R (see Table 2), the other types (C17–C23) are not resolvable by any of the predefined rules. For each conflict, we retrieved and inspected developers’ merged version $m$, using it as the ground truth of conflict resolution.
4.2. Metrics
We defined two evaluation metrics:
Coverage (C) measures among all known conflicts, how many have at least one resolution generated by the resolver:
$C=\dfrac{\text{\# of conflicts with at least one resolution generated}}{\text{Total \# of known conflicts}}$
Accuracy (A) measures among all suggested resolutions, how many of them are correct:
$A=\dfrac{\text{\# of correct resolutions}}{\text{Total \# of generated resolutions}}$
A resolution is correct if the resolved version is semantically equivalent to the ground truth—developers’ resolution recorded in $m$.
4.3. RQ1: BuCoR Resolution Coverage
As shown in Table 4, BuCoR-E and BuCoR-R generated resolutions for 28 and 51 cases, respectively. The coverage metrics they separately achieved are 32% and 58%. When combining their outputs, BuCoR generated at least one resolution for 65 cases, resulting in an overall coverage of 74% (65/88). There are 14 cases where both strategies suggested resolutions. Among the remaining, BuCoR-E and BuCoR-R separately resolved 14 and 37 cases. These numbers imply the two resolution strategies to complement each other.
Listing 2: A conflict without edit example in branches
/* Left Version: PathItem.java */
+ pathItemObject.setRef(apiCallback.callbackUrlExpression());
/* Right Version: Reader.java */
- public void setRef(String ref) {
+ public void set\$ref(String \$ref) {
Listing 3: A conflict whose resolution is unconventional
/* Left Version: JedisCluster.java */
- private int timeout;
/* Right Version: JedisCluster.java */
+ public Set¡String¿ spop(final String key, final long count) {
+   return new JedisClusterCommand¡Set¡String¿¿(connectionHandler, timeout, maxRedirections) {
+    @Override
+   public Set¡String¿ execute(Jedis connection) {
+    return connection.spop(key, count);
+    } }.run(key); }
Listing 4: A conflict whose ground-truth resolution partially overlaps with exemplar edits (haz, 2025)
/* Left Version: */
/* ClientMap.java */
- public final SerializationService getSerializationService() {…}
/* ClientContext.java */
+ public SerializationService getSerializationService() {…}
/* Right Version: ClientMultiMapProxy.java */
+ public boolean put(K key, V value) {
+  Data keyData = getSerializationService().toData(key);
+  Data valueData = getSerializationService().toData(value);
+  PutRequest request = new PutRequest(proxyId, keyData, valueData, -1, ThreadUtil.getThreadId());
+  Boolean result = invoke(request, keyData);
+   return result;
Listing 5: A conflict whose ground-truth resolution has no overlap with exemplar edits (red, 2025)
/* Left Version: RedisClient.java */
+ public RedisClient(String host, int port, int connectTimeout, int commandTimeout) {
+  this(new HashedWheelTimer(), new NioEventLoopGroup(), NioSocketChannel.class, host, port, connectTimeout, commandTimeout);
/* Right Version: RedisClient.java */
- public RedisClient(final Timer timer, EventLoopGroup group, Class¡? extends SocketChannel¿ socketChannelClass, String host, int port, int connectTimeout, int commandTimeout) {
+ public RedisClient(final Timer timer, ExecutorService executor, EventLoopGroup group, Class¡? extends SocketChannel¿ socketChannelClass, String host, int port, int connectTimeout, int commandTimeout) {
Particularly, BuCoR-E suggests nothing when no example exists or no exemplar edit is located in branches. For example, Listing 2 shows a conflict from swagger-core (swa, 2025). The conflict is of type C15—Method: rename def vs. add use, where $r$ renames method setRef(…) and $l$ adds an invocation to that method. Although an intuitive resolution is to update the newly added method call, no edit example in $r$ demonstrates such change. On the other hand, we applied GumTree to compare different versions of source code, and derive AST edit scripts to represent code changes. As GumTree cannot detect any changes to PackageDeclaration or ImportDeclaration, BuCoR-E cannot extract such edits to suggest related resolutions.
BuCoR-R is applicable to many cases where (1) BuCoR-E does not work and (2) the needed resolutions are simple or conventional, such as the case in Listing 2. However, for unconventional conflicts (see C17–C23 in Table 3) that involve entity deletion, parameter addition/deletion, or type change, BuCoR-R generates nothing because there is no typical, generally accepted way to handle them. For example, Listing 3 shows a conflict from Jedis (jed, 2025). This conflict is of C20–Field: remove def vs. add use, where $l$ removes field timeout and $r$ introduces a new reference to timeout. Such a conflict may get resolved by adding back the field def, replacing the added use with something else, or simply removing that use; nevertheless, there is no standard solution or typical resolution pattern followed by developers or documented in literature. Therefore, BuCoR-R did not handle it, while BuCoR-E resolved it in a project-specific way by referring to edit examples. It removed field use as below:
return new JedisClusterCommand<Set<String>>(connectionHandler, maxRedirections) {...}
Among the 21 conflict categories in our dataset, BuCoR resolved conflicts of 19 categories. It did not resolve any conflict of C19 and C22, as (1) BuCoR-E did not find any exemplar edit, and (2) there is no well-accepted resolution pattern for BuCoR to implement.
Finding 1 (Answer to RQ1): BuCoR generated resolution(s) for 65 of the given 88 conflicts, achieving 74% coverage. Among the 79 resolutions it produced, 28 and 51 were separately contributed by BuCoR-E and BuCoR-R.
4.4. RQ2: BuCoR Resolution Accuracy
As shown in Table 4, 21 of the 28 resolutions produced by BuCoR-E are correct, leading to 75% accuracy; 29 of the 51 resolutions output by BuCoR-R are correct, achieving 57% accuracy. Overall, BuCoR obtained 63% accuracy, having 50 of the 79 generated resolutions to be correct. These measurements also show that the two strategies complement each other. Among the 14 cases where both strategies suggest resolutions, BuCoR-E and BuCoR-R correctly resolved 13 and 7 cases, respectively; the correct resolutions overlap in 7 cases. Thus, BuCoR resolved 43 cases correctly.
Particularly, BuCoR-E resolved build conflicts by mimicking developers’ project-specific solutions to related build errors. However, BuCoR-E did not always output correct solutions, because the expected ground-truth resolutions of some conflicts partially overlap with fixes to build errors, or present patterns totally different from those fixes. For instance, in Listing 4, $l$ moves method getSerializationService() from class ClientMap to ClientContext, while $r$ adds two calls to the original method. BuCoR-E replaces the first method call with getContext().getSerializationService(), by following a mined example. Although this replacement is correct, BuCoR-E does not replace the second method call as expected since the original example only contains and updates one method call.
In another scenario (see Listing 5), $r$ updates a constructor’s signature by inserting a parameter of type ExecutorService, and $l$ adds a call to the original constructor. BuCoR-E derived a resolution from branch edits, by updating the call to take an extra argument:
Executors.newFixedThreadPool(Runtime.getRuntime().availableProcessors() * 2).
Although this edit can resolve the conflict, developers’ manual resolution removes the constructor entirely. In such scenarios, branch edits do not reflect developers’ preference of conflict resolution.
Listing 6: A conflict for which the well-accepted resolution pattern does not match developers’ manual resolution (fas, 2025)
/* Left Version: TypeUtils.java */
- import java.beans.Introspector;
/* Right Version: TypeUtils.java */
+ if (compatibleWithJavaBean){
+   propertyName= Introspector.decapitalize(methodName.substring(3));
BuCoR-R sometimes suggested incorrect resolutions, because the frequently applied resolution patterns within it do not always match developers’ practices. For instance, the conflict in Listing 6 has $l$ remove an import and $r$ add a reference to the removed imported class. BuCoR-R resolved this conflict by adding back the removed import. However, developers implemented a replacement method, to avoid the removed dependency. Additionally, for conflicts due to class-hierarchy changes, BuCoR-R modified subclasses to match changes in super classes, while developers sometimes adapted super classes to subclasses.
The 43 conflicts correctly resolved by BuCoR fall into 15 categories, implying that BuCoR can properly resolve diverse conflicts. There are four categories of conflicts for which BuCoR produced some resolutions but none of them are correct: C2, C9, C12, and C18. Three of the categories (C2, C9, and C12) are about mismatches between super types and sub types; two of the categories are caused by parameter list changes (C9 and C18). Among the seven incorrect resolutions suggested by BuCoR-E, two are partially incorrect, by presenting subsets of the needed changes; five are totally irrelevant. Among the 22 incorrect resolutions output by BuCoR-R, 6 are partially incorrect, while 16 are irrelevant.
Finding 2 (Answer to RQ2): BuCoR generated in total 50 correct resolution(s), achieving 63% (50/79) accuracy. BuCoR-E and BuCoR-R separately contributed 21 and 29 correct resolutions.
4.5. RQ3: BuCoR-E vs. BuCoR-R
BuCoR-E and BuCoR-R are different in two aspects:
First, BuCoR-R has higher coverage (58% vs. 32%) by resolving more conflicts; BuCoR-E achieves higher accuracy by having more of its resolutions match developers’ intents (75% vs. 57%). It implies that if developers (1) want to automatically resolve as many conflicts as possible in conventional ways, and (2) have good testing to detect wrong resolutions, they can rely more on BuCoR-R. If developers (1) want to automatically resolve conflicts in unconventional ways but (2) have poor support to detect wrong resolution, they can rely more on BuCoR-E. For instance, when both strategies output resolutions (see Table 1), it is likely that BuCoR-E’s result is better.
Second, the two strategies are good at resolving different types of conflicts. Among the 43 conflicts correctly handled by BuCoR, 7 conflicts are correctly resolved by both strategies. In addition to that, BuCoR-E correctly resolved 14 cases; for 8 of these cases, BuCoR-R could not suggest anything. The eight cases are related to entity removal (i.e., C17, C20, C23) and parameter-list changes (i.e., C21). It means that BuCoR-E is better at resolving conflicts in unconventional or complex ways. Meanwhile, BuCoR-R correctly resolved 22 cases, for which BuCoR-E suggested nothing. These 22 conflicts are mainly related to super-sub mismatches (C3–C4, C8, C10–C11) and duplicated entities (C14, C16). It means that BuCoR-R is better at handling simple conflicts in conventional ways.
4.6. Runtime Overhead
On our dataset, BuCoR spent 0.1–53.6 minutes on each of the 55 merging scenarios, with 2.9 minutes as the mean and 0.8 minutes as the median. Among the four components of BuCoR-E, conflict detection and graph construction are the most time-consuming, roughly taking up 55% and 36% of the overall runtime. The time BuCoR spent on each merging scenario is closely related to the number of (1) Java files and (2) program entities. Namely, the more files to parse and the more entities to analyze, the longer BuCoR takes. We noticed that BuCoR spent the least time 0.1 minutes, when there are 20 Java files and 1,229 program entities under processing. It spent the most time 53.6 minutes when there are 1,472 Java files and 48,299 entities being analyzed.
5. Threats to Validity
Threats to External Validity. Our evaluation dataset consists of 88 real-world build conflicts, spanning 21 distinct conflict types. However, the program contexts and conflict patterns covered by this dataset may not fully capture the diversity of build conflicts in real world. BuCoR-R includes 16 well-defined rules of handling frequently occurring conflicts. Although they cover all the typical conflict-resolution strategies we are aware of, they may not include all the frequently applied strategies in reality. In the future, to enhance our research representativeness, we will (1) expand our dataset to include more merging scenarios, and (2) enlarge the ruleset of BuCoR-R whenever possible.
Threats to Internal Validity. BuCoR-E leverages the threshold 0.618 to decide whether two AST nodes are similar enough to have a value match, because prior work (Shen et al., 2019) shows that this setting led to reasonably good results. However, there can be scenarios where two matching AST nodes have a lower similarity score than 0.618, disabling BuCoR-E to apply edits as expected. In the future, we will explore better ways of node matching.
Threats to Construct Validity. BuCoR leverages Bucond to detect conflicts. If Bucond cannot detect certain build conflicts (e.g., conflicts in build scripts), BuCoR cannot resolve those conflicts either. Prior work (Towqir et al., 2023) shows that Bucond has a very good coverage of conflict types, and there is a low chance of Bucond to miss real-world conflicts in Java source code. Therefore, the effectiveness of BuCoR is not considerably limited by Bucond.
6. Related Work
The related work of our research includes automated software merge, and empirical studies on merge conflicts.
6.1. Automated Software Merge
Software Merge Tools (Apel et al., 2011; Brun et al., 2011; Apel et al., 2012; Guimarães and Silva, 2012; Zhu and He, 2018; Shen et al., 2019). FSTMerge (Apel et al., 2011; Cavalcanti et al., 2017; jFS, 2021) parses Java code for ASTs, and matches nodes between branches based on the class or method signatures; it then integrates the edits inside each matching pair via textual merge. JDime (Apel et al., 2012) also parses Java code for ASTs; however, unlike FSTMerge, it merges code purely via structural matching between ASTs and tree manipulation. AutoMerge (Zhu and He, 2018) extends JDime, by proposing multiple alternative strategies to resolve conflicts between branches, with each strategy integrating branch edits in a unique way. IntelliMerge (Shen et al., 2019) presents a graph-based refactoring-aware merging algorithm. These tools can resolve some textual conflicts, but cannot handle build conflicts. Crystal (Brun et al., 2011) and WeCode (Guimarães and Silva, 2012) help reveal three kinds of conflicts. Both tools first apply textual merge to create a merged software, and reveal textual conflicts along the way. They then adopt automatic build and testing to find build/test errors, regarding those errors as indicators of build and test conflicts. However, neither tool pinpoints or resolves build/test conflicts.
Detectors of Build Conflicts (Wuensche et al., 2020b; Towqir et al., 2023). The existing tools both statically analyze branch versions, reason about branch edits, and contrast extracted edits with predefined patterns to detect conflicts. However, neither tool automatically resolves build conflicts.
Detectors of Test Conflicts (Sousa et al., 2018; Da Silva et al., 2024). SafeMerge (Sousa et al., 2018) takes in $b$, $l$, $r$, and $m$, for a given merging scenario. It statically infers the relational postconditions of distinct versions to model program semantics. By comparing postconditions, SafeMerge decides whether $m$ is free of conflicts, i.e., without introducing new semantics nonexistent in $l$ or $r$. SAM (Da Silva et al., 2024) also takes in four program versions related to a merging scenario. It randomly generates tests with EvoSuite, to explore situations where $b$, $l$, $r$ pass a test but $m$ fails that test.
Resolution of Textual Conflicts (Pan et al., 2021; Aldndni et al., 2023, 2024; Elias et al., 2023; Dinella et al., 2021; Svyatkovskiy et al., 2022). RPredictor (Aldndni et al., 2023, 2024) and MESTRE (Elias et al., 2023) train machine learning (ML) models with traditional algorithms (e.g., Random Forest), using features to characterize conflict content, merging scenarios, evolution history, developer experience, and/or edits applied by different branches. Given a conflict, the models then predict what resolution strategy to apply (e.g., keep left). DeepMerge (Dinella et al., 2021) and MergeBERT (Svyatkovskiy et al., 2022) employ deep-learning to resolve specified conflicting chunks automatically. Pan et al. (Pan et al., 2021) defined a domain-specific language (DSL) to capture repetitive resolution patterns, and proposed a program synthesis approach to learn resolution strategies as DSL programs from example resolutions.
Resolvers of Build Conflicts (Sung et al., 2020; Zhang et al., 2022). MrgBldBrkFixer (Sung et al., 2020) compares the ASTs of C++ files, to resolve conflicts related to renaming changes. Gmerge (Zhang et al., 2022) applies few-shot learning to GPT-3, to resolve conflicts due to renaming. They are analogous to BuCoR-E, because they all infer and apply patterns by analyzing exemplar edits. We did not empirically compare BuCoR with either tool, as they are close-sourced tools to handle conflicts in C/C++ code.
BuCoR is more advanced in three aspects. First, it defines a comprehensive set of 16 resolution rules: in addition to conflicts due to renaming changes, these rules also handle conflicts due to class-hierarchy change and entity addition/deletion. Second, it defines sophisticated algorithms to (1) extract resolution-related edits from branches via control- and data- dependency analysis, (2) match patterns with program context in a use-centric manner, and (3) derive candidate resolutions to the same conflict by tentatively matching alternative patterns with the same edit context. Thus, in addition to symbol renaming, it resolves conflicts through systematic code editing to consistently revise program semantics. Third, it defines a hybrid approach between example-based and rule-based resolutions, to combine the advantages of both methodologies.
6.2. Empirical Studies of Merge Conflicts
Relationship between Merge Conflicts and Software Maintenance (Estler et al., 2014; Ahmed et al., 2017; Mahmoudi et al., 2019). Estler et al. (Estler et al., 2014) surveyed 105 student developers, to study the relationship between awareness (i.e., knowing “who’s changing what”) and merge conflicts. Ahmed et al. (Ahmed et al., 2017) explored the effect of bad design (code smells) on merge conflicts. Mahmoudi et al. (Mahmoudi et al., 2019) studied the relation between merge conflicts and 15 popular refactoring types. These researchers showed that merge conflicts are widespread, although they studied textual conflicts.
Characterization of Merge Conflicts (Ghiotto et al., 2018; Shen et al., 2022; Shen and Meng, 2024; Da Silva et al., 2022). Ghiotto et al. (Ghiotto et al., 2018) and Shen et al. (Shen and Meng, 2024) characterized textual conflicts from open-source Java projects in terms of number of chunks, size, program constructs involved, their validity (i.e., true vs. false positives), edit types, file types, and/or manual resolution strategies. However, neither work studies build conflicts.
Shen et al. (Shen et al., 2022) followed the methodology of Crystal (Brun et al., 2011) and WeCode (Guimarães and Silva, 2012) to reveal 3 types of conflicts in 208 open-source repositories. They manually inspected in total 538 conflicts to characterize the root causes and developers’ resolution strategies of all conflict types. Da Silva et al. (Da Silva et al., 2022) collected 239 build conflicts to study their root causes and resolution patterns. Both studies reported that most build conflicts are caused by declarations removed or updated by one branch but referenced by another branch. Da Silva et al.  also observed that conflicts caused by renaming are often resolved by updating the missing reference, whereas removed declarations are often reintroduced. Both studies motivated our research, and present the initial datasets for us to use when creating our own evaluation dataset.
7. Conclusion
This paper introduces BuCoR, a novel conflict resolver to conduct example-based and rule-based transformation. It applies program static analysis to (1) detect conflicts, (2) mine edit examples in branches, (3) derive transformation patterns from mined examples, (4) and apply inferred or predefined patterns to resolve conflicts. Compared with prior work, it significantly pushes the boundary of automatic conflict resolution. Our investigation is deeper as (a) the conflicts we focus on are very diverse; (b) the edits we suggest vary a lot depending on the program context and conflict types; (c) in the scenarios where no standard resolution pattern exists, BuCoR can create resolutions by locating, refining, and reusing relevant edits. In the future, we will improve BuCoR by (i) expanding the rule set of BuCoR-R, and (ii) improving the example mining capability of BuCoR-E when no local example is extractable from branches.

### v2

@@
-Resolving Build Conflicts via Example-Based and Rule-Based Program Transformations
+Combining Example-Based and Rule-Based Program Transformations to Resolve Build Conflicts
 Abstract.
@@
 They degrade software quality and hinder programmer productivity.
-While several tools detect build conflicts, few offer meaningful support for resolvinntional cases like those caused by method removal.
+While several tools detect build conflicts, few offer meaningful support for resolving them.
 To overcome limitations of existing tools, we introduce BuCoR (BUild COnflict Resolver), a new conflict resolver.
@@
 To resolve conflicts, it employs two complementary strategies: example-based transformation (BuCoR-E) and rule-based transformation (BuCoR-R).
-BuCoR-R applies predefined rules to handle common, well-understood conflicts.
+BuCoR-R applies predefined rules to resolve conflicts in frequently suggested or conventional ways.
 BuCoR-E mines branch versions ($l$ and $r$) for exemplar edits applied to fix related build errors.
-From these examples, it infers and generalizes program transformation patterns to resolve more complex conflicts.
+From these examples, it infers and generalizes program transformation patterns to resolve conflicts in project-specific or unconventional ways.
 We evaluated BuCoR on 88 real-world build conflicts spanning 21 distinct conflict types.
-BuCoR generated at least one solution for 65 cases, and correctly resolved 43 conflicts.
+BuCoR generated at least one solution for 65 cases, and correctly resolved 34 conflicts.
 We observed that this hybrid approach—combining context-aware, example-based learning with structured, rule-based resolution—can effectively help resolve conflicts.
@@
 software merge, build conflict, static analysis, example-based, rule-based, program transformation, example mining, pattern inference
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software evolution
+journalyear: 2026copyright: ccconference: 23rd International Conference on Mining Software Repositories; April 13–14, 2026; Rio de Janeiro, Brazilbooktitle: 23rd International Conference on Mining Software Repositories (MSR ’26), April 13–14, 2026, Rio de Janeiro, Brazildoi: 10.1145/3793302.3793347isbn: 979-8-4007-2474-9/2026/04ccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software evolution
 1. Introduction
@@
 Prior work shows that developers often spend hours or days detecting and resolving conflicts before correctly merging branches (Kasi and Sarma, 2013).
-As illustrated in Figure 1, a typical merging scenario involves five program versions: two branch versions $l$ and $r$ whose edits need to be merged, base version $b$—the common origin of both branches, the automatically merged version $A_{m}$ produced by git-merge when it naïvely integrates branch edits textually, and the manually merged version $m$ that developers create based on $A_{m}$.
+As illustrated in Figure 1, a typical merging scenario involves five program versions: two branch versions $l$ and $r$ whose edits need to be merged, base version bb—the common origin of both branches, the automatically merged version $A_{m}$ produced by git-merge when it naïvely integrates branch edits textually, and the manually merged version $m$ that developers create based on $A_{m}$.
 Among the various possible conflicts between $l$ and $r$, build conflicts refer to the incompatible edits whose naïve integration triggers build errors in the resulting merged version.
@@
 Namely, the newly added method call $m()$ is not associated with any defined method.
-Figure 1. A typical merging scenario can involve up to five program versions
-Although tools were created to detect build conflicts (Towqir et al., 2023; Wuensche et al., 2020a; Shen et al., 2019), there is rare automatic support to resolve such conflicts.
-In particular, Gmerge (Zhang et al., 2022) relies on Clang compiler message to locate conflict-related changes in branches, and applies k-shot learning with GPT-3 to suggest symbol renaming for conflict resolution.
+Figure 1. A merging scenario involves five program versions
+Figure 2. A build conflict due to the updated def of m() by $r$, and added use by $l$
+Figure 3. BuCoR leverages Bucond (Towqir et al., 2023) to detect build conflicts, and employs two complementary strategies to resolve conflicts
+Although tools were created to detect build conflicts (Towqir et al., 2023; Wuensche et al., 2020), there is rare automatic support to resolve such conflicts.
+In particular, Gmerge (Zhang et al., 2022) relies on Clang compiler messages to locate conflict-related changes in branches, and applies k-shot learning with GPT-3 to suggest symbol renaming for conflict resolution.
 MrgBldBrkFixer (Sung et al., 2020) starts with a naïvely merged version $A_{m}$, looks for build errors due to failed resolution of symbols, identifies symbol renaming changes in branches-to-merge, extracts related patches in either branch, and similarly applies those patches to fix build conflicts.
-These tools suffer from three limitations:
-Figure 2. A build conflict due to the updated def of m() by $r$, and added use by $l$
-Figure 3. BuCoR leverages Bucond (Towqir et al., 2023) to detect build conflicts, and employs two complimentary strategies to resolve conflicts
-  1. (1)
-    Both tools adopt compiler errors to locate conflicts; if compilers fail to locate errors, neither tool works.
+These tools suffer from two limitations:
   1. (1)
@@
     They suggest simple edits of symbol renaming to adapt uses of renamed entities, but fail to suggest complex edits that systematically change uses and surrounding context.
-To overcome those limitations, we introduce BuCoR (BUild COnflict Resolver), a novel resolver of build conflicts.
-As shown in Figure 3, given the three program versions ($l$, $b$, $r$) related to a merging scenario, BuCoR first applies Bucond (Towqir et al., 2023) to reveal conflicting edits between branches.
-As Bucond detects conflicts via static analysis instead of automatic build, BuCoR eliminates dependencies on compilers.
+To overcome these limitations, we introduce BuCoR (BUild COnflict Resolver), a novel resolver of build conflicts.
+As shown in Figure 3, given the three program versions ($l$, $b$, $r$) related to a merging scenario, BuCoR first applies Bucond (Towqir et al., 2023), a static analysis-based tool, to reveal conflicting edits between branches.
 BuCoR also adopts git-merge to generate a naïvely merged version $A_{m}$, extending Bucond to (1) create a graphical representation for $A_{m}$ and (2) relate that graph with those of the three input program versions.
@@
 If a full or partial match is found, BuCoR-E customizes and applies the entire or partial pattern for conflict resolution.
-Rule-based transformation (BuCoR-R): Prior studies reveal patterns frequently applied to resolve certain kinds of build conflicts (Shen et al., 2022; Da Silva et al., 2022).
+Rule-based transformation (BuCoR-R): Prior studies reveal patterns frequently applied to resolve certain kinds of build conflicts (Wuensche et al., 2020; Sung et al., 2020).
 Following those studies, we defined and implemented 16 resolution rules/patterns in BuCoR-R.
@@
 To evaluate the tool effectiveness, we applied BuCoR to 88 real-world build conflicts in 30 open-source projects.
-BuCoR correctly resolved 43 conflicts.
-BuCoR-E and BuCoR-R separately generated resolutions for 28 and 51 conflicts; 21 and 29 of those resolutions separately match developers’ resolutions recorded in version history.
+BuCoR correctly resolved 34 conflicts.
+BuCoR-E and BuCoR-R separately generated resolutions for 28 and 51 conflicts; 21 and 20 of those resolutions separately match developers’ resolutions recorded in version history.
 All these numbers demonstrate BuCoR’s effectiveness in handling conflicts.
@@
   • We created BuCoR, a new static analysis-based resolver of build conflicts, to novelly combine example-based transformation with rule-based transformation.
-Unlike prior work, BuCoR does not adopt compilers to detect conflicts.
   • We explored BuCoR-E, an advanced example-based approach of conflict resolution.
@@
   • We systematically evaluated BuCoR with a dataset of 88 conflicts that span 21 types, and observed novel phenomena.
+At https://github.com/shadabtowqir/BuCoR, we open-sourced our program and data.
 Table 1. A motivating example of build conflict, whose resolution requires context-specific edits more than symbol renaming
@@
 Changes in $\boldsymbol{l}$ (responsible def updates): In TypeSerializerConfig.java, names of the file, Java class, and constructor are all updated to SerializerConfig.
-Changes in $\boldsymbol{r}$ (responsible use introduction): In a newly added file XmlClientConfigBuilder.java, a method is defined to access TypeSerializerConfig + private void handleSerializers(Node node, SerializationConfig serializationConfig) { + + if (”type-serializer”.equals(name)) { + TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); + typeSerializerConfig.setClassName(value); + final String typeClassName = getAttribute(child, ”type-class”); + typeSerializerConfig.setTypeClassName(typeClassName); + serializationConfig.addTypeSerializer(typeSerializerConfig); + } } Edit example in $\boldsymbol{l}$ to adapt usage of TypeSerializerConfig: private void handleSerializers(Node node, …) { - if(”type-serializer”.equals(name)){ - TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); - typeSerializerConfig.setClassName(value); + if(”serializer”.equals(name)){ + SerializerConfig serializerConfig = new SerializerConfig(); + serializerConfig.setClassName(value); final String typeClassName=retrieveAttribute(child, ”type-class”); - typeSerializerConfig.setTypeClassName(typeClassName); - serializationConfig.addTypeSerializer(typeSerializerConfig); + serializerConfig.setTypeClassName(typeClassName); + serializationConfig.addSerializerConfig(serializerConfig);
-Tool-generated edits that are applicable to $A_{m}$, to resolve conflicts
-(c) A naïve resolution producible by existing tools and BuCoR-R (d) A more comprehensive resolution produced by BuCoR-E
-private void handleSerializers(Node node, …) { if (”type-serializer”.equals(name)) { -   TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); +   SerializerConfig typeSerializerConfig = new SerializerConfig(); typeSerializerConfig.setClassName(value); final String typeClassName=getAttribute(child, ”type-class”); typeSerializerConfig.setTypeClassName(typeClassName); serializationConfig.addTypeSerializer(typeSerializerConfig); } } private void handleSerializers(Node node, …) { - if (”type-serializer”.equals(name2)) { -  TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); -  typeSerializerConfig.setClassName(value); + if (”serializer”.equals(name2)) { +  SerializerConfig serializerConfig = new SerializerConfig(); +  serializerConfig.setClassName(value); final String typeClassName=getAttribute(child, ”type-class”); -  typeSerializerConfig.setTypeClassName(typeClassName); -  serializationConfig.addTypeSerializer(typeSerializerConfig); +  serializerConfig.setTypeClassName(typeClassName); +  serializationConfig.addSerializerConfig(serializerConfig); } }
+Changes in $\boldsymbol{r}$ (responsible use introduction): In a newly added file XmlClientConfigBuilder.java, a method is defined to access TypeSerializerConfig + private void handleSerializers(Node node, SerializationConfig serializationConfig) { + + if (”type-serializer”.equals(name)) { + TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); + typeSerializerConfig.setClassName(value); + final String typeClassName = getAttribute(child, ”type-class”); + typeSerializerConfig.setTypeClassName(typeClassName); + serializationConfig.addTypeSerializer(typeSerializerConfig); + } }      Edit example in $\boldsymbol{l}$ to adapt usage of TypeSerializerConfig: private void handleSerializers(Node node, …) { - if(”type-serializer”.equals(name)){ - TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); - typeSerializerConfig.setClassName(value); + if(”serializer”.equals(name)){ + SerializerConfig serializerConfig = new SerializerConfig(); + serializerConfig.setClassName(value); final String typeClassName=retrieveAttribute(child, ”type-class”); - typeSerializerConfig.setTypeClassName(typeClassName); - serializationConfig.addTypeSerializer(typeSerializerConfig); + serializerConfig.setTypeClassName(typeClassName); + serializationConfig.addSerializerConfig(serializerConfig);    Tool-generated edits that are applicable to $A_{m}$, to resolve conflicts    (c) A naïve resolution producible by existing tools and BuCoR-R (d) A more comprehensive resolution produced by BuCoR-E private void handleSerializers(Node node, …) { if (”type-serializer”.equals(name)) { - TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); + SerializerConfig typeSerializerConfig = new SerializerConfig(); typeSerializerConfig.setClassName(value); final String typeClassName=getAttribute(child, ”type-class”); typeSerializerConfig.setTypeClassName(typeClassName); serializationConfig.addTypeSerializer(typeSerializerConfig); } } - if (”type-serializer”.equals(name2)) { - TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); - typeSerializerConfig.setClassName(value); + if (”serializer”.equals(name2)) { + SerializerConfig serializerConfig = new SerializerConfig(); + serializerConfig.setClassName(value); final String typeClassName=getAttribute(child, ”type-class”); - typeSerializerConfig.setTypeClassName(typeClassName); - serializationConfig.addTypeSerializer(typeSerializerConfig); + serializerConfig.setTypeClassName(typeClassName); + serializationConfig.addSerializerConfig(serializerConfig); } }
 2. A Motivating Example
-To facilitate discussion, this section introduces a running example we crafted based on a real-world project hazelcast (haz, 2022).
-As shown in Table 1(a), a merging scenario has $l$ rename a class and $r$ insert code to use the original class.
+To facilitate discussion, we introduce a running example we crafted based on a real-world project hazelcast (haz, 2022).
+As shown in Table 1(a), a merging scenario has $l$ rename a class TypeSerializerConfig and $r$ insert code to use the original class.
 The naïve integration of these edits can trigger a build error as the newly introduced uses refer to a nonexistent class def; thus, a build conflict occurs.
-To resolve such a conflict, existing tools simply update class references by replacing TypeSerializerConfig with SerializerConfig (Table 1(c)).
-However, such a replacement is insufficient, as the context still has variable typeSerializerConfig and literal “type-serializer” match the original class usage.
+To resolve such conflicts, existing tools update class references by replacing TypeSerializerConfig with SerializerConfig (Table 1(c)).
+However, such a replacement is insufficient, as the context still has variable typeSerializerConfig and literal “type-serializer” match the original class.
 Consequently, after existing tools update class references, developers need to manually replace defs and uses of related variables/literals, to ensure consistent updates and prevent semantic conflicts.
-Furthermore, when such class uses are introduced at multiple places, developers have to go over all places to manually apply those edits again and again, which process is tedious and error-prone.
+When such class uses are introduced at multiple places, developers have to go over all places to manually apply those edits again and again, which process is tedious and error-prone.
 To overcome the limitations of existing tools and better help developers, we introduce BuCoR—a hybrid approach to combine example-based resolution with rule-based resolution.
 The example-based resolution BuCoR-E is derived from our insight, on the association between build errors in branches and build conflicts in the merged version.
-Basically, many build errors are caused by mismatches between def and use of the same program entities; many build conflicts are caused by mismatches between modified def and newly introduced use of entities (Shen et al., 2022; Da Silva et al., 2022).
-If on either branch, developers resolved def-use mismatches by applying specialized edits; then they are likely to reapply the same or similar edits to resolve conflicts that show the same kind of mismatches in software merge.
+Basically, many build errors are caused by mismatches between def and use of the same program entities.
+There is commonality between (1) fixes to build errors on branches, and (2) resolutions to build conflicts in the merged version  (Shen et al., 2022; Da Silva et al., 2022).
+If on either branch, developers resolved def-use mismatches by applying specialized edits; then they are likely to reapply similar edits to resolve conflicts that show the same kind of mismatches in software merge.
 For each reported conflict, BuCoR-E locates the responsible def-change, and mines the contributing branch for developers’ exemplar edit $E$ applied to adjust existing uses of the changed entity.
@@
 BuCoR-E then derives a transformation pattern $P$ from $E$, by abstracting away irrelevant edit detail and/or program context.
-For the responsible use-introduction of each conflict, BuCoR-E tentatively establishes context matching between $P$ and the edit location.
-If the matching succeeds, BuCoR-E customizes $P$ by generating edit operations with respect to the matched context.
+For the responsible use-introduction of each conflict, BuCoR-E establishes context matching between $P$ and the edit location.
+If a match is found, BuCoR-E customizes $P$ by creating edit operations with respect to the matched context.
 BuCoR then applies those operations to transform code.
 As shown in Table 1(d), the context-to-handle is different from the original edit context (see Table 1(b)): it uses a different variable (name2 vs.
-name), and invokes a different method (getAttribute(...) vs.
+name), and calls a different method (getAttribute(...) vs.
 retrieveAttribute(...)).
-Albeit the differences, BuCoR still resolves the conflict by mimicking developers’ coding practices.
+However, BuCoR could resolve the conflict by mimicking developers’ edits.
 In addition to BuCoR-E, our tool also integrates a rule-based transformation approach BuCoR-R.
@@
 3.1. Conflict Detection (Bucond (Towqir et al., 2023))
-[…]
-For implementation, Bucond uses JavaParser to parse source code, and adopts JGraphT (jgr, 2025) to construct and analyze PEGs.
+Bucond statically analyzes three program versions related to each merging scenario—$b$, $l$, rr—to identify build conflicts.
+It models each version as a program entity graph (PEG), which captures defined program entities (e.g., classes and methods) and their relations (e.g., type reference or method calls).
+By comparing PEGs, Bucond extracts entity-level edits in $l$ and $r$.
+It then matches these edits with 57 predefined patterns of conflicting edits.
+For instance, one pattern checks when a method is renamed in one branch, whether the other branch adds any call to the original method.
+Bucond reports a conflict if any match is found.
+Unlike compiler-based tools, Bucond does not try to generate or build the naïvely merged version $A_{m}$.
+It can detect conflicts even if $A_{m}$ is uncompilable, and pinpoint the specific edits responsible—which compilers cannot do.
+For implementation, Bucond uses JavaParser (jp, 2025) to parse source code, and adopts JGraphT (jgr, 2025) to construct and analyze PEGs.
 3.2. Graph Construction
@@
 3.3.1. Phase I. Example Mining
-Given a conflict, BuCoR-E identifies the responsible def-change, mines the branch contributing that change, and looks for any exemplar edit which adapts entity uses and surrounding code for that change.
-In our motivating example, because $l$ renames class TypeSerializerConfig and its constructor, it is the contributor branch of def-change.
-By traversing entity-level edits of $l$$\Delta(G_{l},G_{b})$, BuCoR-E locates changes of the corresponding class/constructor references.
-Namely, if an entity $\boldsymbol{e}$ uses TypeSerializerConfig in $b$ but uses SerializerConfig in $l$, then the entity $e$ contains an edit example.
-Certainly, if multiple entities adapt their usage to the same def-change, BuCoR-E considers all these entities to have relevant examples.
+[…]
 To extract and represent each edit example, BuCoR-E applies an off-the-shelf syntactic differencing tool GumTree (Falleri et al., 2014) to the base and branch versions of $e$, to generate an edit script of Abstract Syntax Tree (AST).
 The script may have four kinds of operations:
-[…]
-[…]
-[…]
-[…]
+  • update($t$, $v_{n}$): To replace the old value of node $t$ with the new value $v_{n}$.
+  • add($t$, $t_{p}$, $i$, $l$, $v$): To add a new node $t$ to the AST, as the $i^{th}$ child of node $t_{p}$.
+Here, $l$ and $v$ separately specify $t$’s entity type (e.g., method invocation) and its value (e.g., the statement string).
+  • delete($t$): To remove a node $t$ from the AST.
+  • move($t$, $t_{p}$, $i$): To move a node $t$ and its subtree to the $i^{th}$ child of node $t_{p}$.
 For our motivating example, the script extracted from the located exemplar edit includes eight operations: each operation updates a node to replace a string literal, type usage, constructor/method usage, or variable usage (see Figure 5).
@@
 Output: Node matches between the ASTs: $W$
-/* Find a match for the last usage of $e$/
+/* Find a match for the last usage of $e$ */
 0.1 Locate a statement $s_{p}\in P$, which has the last usage of $e$
@@
 0.18
- /* Compare parent nodes of $s_{p}$ and $s_{m}$/
+ /* Compare parent nodes of $s_{p}$ and $s_{m}$ */
 0.19  Locate the parent node of $s_{p}$, $p_{p}\in P$, if there is any
@@
 We designed an algorithm to perform use-centric, bottom-up context matching.
-As illustrated by Algorithm 1, given a pattern $P$, a method $M$ whose entity use causes a conflict, and the changed entity $e$, BuCoR-E tentatively matches nodes in a bottom-up manner with the best effort.
-Specifically, it first locates a statement $s_{p}$ in $P$ that contains the last usage of $e$, to search among $M$ for a statement $s_{m}$ best matching $s_{p}$ (see 1.1–1.2).
+As shown in Algorithm 1, given a pattern $P$, a method $M$ whose entity use causes a conflict, and the changed entity $e$, BuCoR-E performs best-effort bottom-up node matching.
+It first locates in $P$ a statement $s_{p}$ that contains the last usage of $e$, and searches $M$ for a statement $s_{m}$ best matching $s_{p}$ (lines 1.1–1.2).
 To compare statements, we check their AST node types and values to calculate a matching score as below:
@@
 Type match enables patterns to get applied to the context (1) distinct from the original edit example, but (2) preserving the structural consistency.
-  • Value Match: Two nodes, $x$ and $y$, have a value match if their strings have a 3-gram similarity score greater than 0.618.
-Here, the string of a simple statement (e.g, MethodInvocation) includes all content of that statement; the string of a complex statement (e.g., conditional or loop structure) only includes the structure header (e.g., if-condition or for-loop header).
-A successful value match increments the overall matching score by the calculated similarity value.
-This threshold-based approach allows similar nodes to match, even though they are not identical.
-Such a flexibility also increases the applicability of inferred patterns.
+  • Value Match: Two nodes, $x$ and $y$, have a value match if their strings have a 3-gram similarity score greater than 0.618; the similarity score is added to the total matching score.
+The string of a simple statement (e.g., MethodInvocation) includes all content of that statement; the string of a complex statement (e.g., loop) only includes the header.
+This threshold-based approach allows flexible matching across similar code, and increases the applicability of inferred patterns.
 Particularly, if $x$ and $y$ are identical, they have a type match and the 3-gram similarity is 1, so the overall matching score is 2.
-Once a best match of $s_{p}$ is found (with score $>$ 1.618), BuCoR-E treats the pair $(s_{p},s_{m})$ as anchor points to guide node matching for their siblings (see 1.5–1.14).
-Specifically for siblings preceding $s_{p}$, it enumerates nodes in reverse order: it first compares $s_{p}$’s nearest sibling against all siblings preceding $s_{m}$ for a best match; it then proceeds to match $s_{p}$’s more distant siblings against any remaining unmatched nodes, moving backward from the most recently matched one.
-Similarly, for siblings following $s_{p}$, BuCoR-E matches nodes in forward order.
-After searching best matches for all sibling nodes of $s_{p}$, the process continues to tentatively match the parent node of $s_{p}$ with that of $s_{m}$.
+Once a best match of $s_{p}$ is found (with score >> 1.618), BuCoR-E treats the pair $(s_{p},s_{m})$ as anchor points to guide sibling match (lines 1.5–1.14).
+Preceding siblings are matched in reverse order, starting from the nearest sibling, while following siblings are matched in forward order.
+The 1.618 threshold reflects the expectation of both a type match and a value similarity of at least 0.618.
+Here, 0.618 is the golden ratio used in prior work (Shen et al., 2019) to yield robust matching.
+After matching all siblings, BuCoR-E attempts to match the parent nodes of $s_{p}$ and $s_{m}$.
 If this match succeeds, BuCoR-E repeats 1.3–1.17 in Algorithm 1, using the newly matched blocks as anchors for further match.
-The procedure continues until no more parent node in $P$ can be found or the parent match fails.
+The procedure continues until no further parent matches are possible.
 In the end, the algorithm outputs best matches of nodes, and their matching scores.
-If Phase II infers multiple patterns from edit examples, these patterns may be all applicable to the method-to-edit $M$.
-To find the best one, we defined two ways to score the overall context matching between each pattern and $M$, based on the output of Algorithm 1:
+If multiple patterns inferred from Phase II apply to method $M$, BuCoR-E ranks them with two metrics derived from Algorithm 1:
   • Sum of Matching Scores ($\Sigma_{m}$): Adding up the scores of all best matches.
@@
 Thus, among all applicable patterns, we select the one with the highest $\Sigma_{m}$.
-If multiple candidates have a tie, we choose the one with the highest $C_{m}$, ensuring a closer semantic match between that pattern and $M$.
+Ties are broken using $C_{m}$ to favor closer semantic matches.
 In our motivating example, the method-to-edit (Table 1(c)) shares a highly similar context with the inferred pattern (Listing 1): $\Sigma_{m}=9.94$ and $C_{m}=4$.
-All five edited statements in the pattern sequentially find best matches within the method: four statements match exactly, and the if-condition has a highly similar counterpart.
+All five edited statements in the pattern sequentially find best matches within the method: with four exact matches and a highly similar if-condition.
 Step 2 of Phase III: AST Rewriting.
@@
 BuCoR-R resolves conflicts via rule-based program transformations.
-To define practical resolution rules or patterns, we manually inspected the rules and data observed by prior empirical studies (Shen et al., 2019; Da Silva et al., 2022), and expanded those rule sets to generalize rules across similar but different conflict types.
-Table 2 shows the resulting 16 conventional patterns we implemented in BuCoR-R; each pattern resolves one type of frequent conflicts.
-Among the 16 conflict types, only 6 types involve symbol renaming (C1, C6, C7, C11, C13, C15).
+To define practical resolution rules or patterns, we manually inspected the rules and data observed by prior empirical studies (Sung et al., 2020; Da Silva et al., 2022), and expanded those rule sets to generalize rules across similar but different conflict types.
+Table 2 shows the resulting 16 conventional rules we implemented in BuCoR-R; each rule resolves one type of frequent conflicts.
+Among the 16 rules, only 6 rules involve symbol renaming (C1, C6, C7, C11, C13, C15).
 As prior work only resolves conflicts caused by symbol renaming, they cannot handle 10 of the conflict types BuCoR-R focuses on.
+Only 8 of the 16 rules are clearly described by prior work: C1, C3, C5–C6, C8, C14–C16.
+We derived the remaining rules by generalizing across related conflict types.
 Table 2. The 16 resolution rules used in BuCoR-R
@@
 BuCoR-E 28 21 32% (28/88) 75% (21/28)
-BuCoR-R 51 29 58% (51/88) 57% (29/51)
-BuCoR 79 50 74% (65/88) 63% (50/79)
+BuCoR-R 51 20 58% (51/88) 39% (20/51)
+BuCoR 79 41 74% (65/88) 52% (41/79)
 4. Experiment
@@
 4.1. Dataset
-We constructed our evaluation dataset by reusing the datasets of prior work (Towqir et al., 2023; Da Silva et al., 2022; Shen et al., 2019) with our best effort.
+We constructed our evaluation dataset by reusing the datasets of prior work (Towqir et al., 2023; Da Silva et al., 2022) with our best effort.
 Given a merging scenario with at least one known build conflict, we decided whether it is reusable based on the following criteria:
@@
 Coverage (C) measures among all known conflicts, how many have at least one resolution generated by the resolver:
-$C=\dfrac{\text{\# of conflicts with at least one resolution generated}}{\text{Total \# of known conflicts}}$
+[…]
 Accuracy (A) measures among all suggested resolutions, how many of them are correct:
@@
 A resolution is correct if the resolved version is semantically equivalent to the ground truth—developers’ resolution recorded in $m$.
+Two authors independently inspected BuCoR’s resolution to examine semantic alignment, and had a discussion whenever they disagree.
 4.3. RQ1: BuCoR Resolution Coverage
@@
 Listing 2: A conflict without edit example in branches
-/* Left Version: PathItem.java */
+/* Left Version: PathItem.java adds a method call to setRef(…) */
 + pathItemObject.setRef(apiCallback.callbackUrlExpression());
-/* Right Version: Reader.java */
+/* Right Version: Reader.java renames setRef(…) */
 - public void setRef(String ref) {
@@
 Listing 3: A conflict whose resolution is unconventional
-/* Left Version: JedisCluster.java */
+/* Left Version: JedisCluster.java removes a field */
 - private int timeout;
-/* Right Version: JedisCluster.java */
+/* Right Version: JedisCluster.java adds a reference to field timeout */
 + public Set¡String¿ spop(final String key, final long count) {
@@
 Listing 4: A conflict whose ground-truth resolution partially overlaps with exemplar edits (haz, 2025)
-/* Left Version: */
+/* Left Version moves a method between classes */
 /* ClientMap.java */
@@
 + public SerializationService getSerializationService() {…}
-/* Right Version: ClientMultiMapProxy.java */
+/* Right Version: ClientMultiMapProxy.java adds calls to the original method */
 + public boolean put(K key, V value) {
@@
 Listing 5: A conflict whose ground-truth resolution has no overlap with exemplar edits (red, 2025)
-/* Left Version: RedisClient.java */
+/* Left Version: RedisClient.java adds a call to one of its constructors */
 + public RedisClient(String host, int port, int connectTimeout, int commandTimeout) {
 + this(new HashedWheelTimer(), new NioEventLoopGroup(), NioSocketChannel.class, host, port, connectTimeout, commandTimeout);
-/* Right Version: RedisClient.java */
+/* Right Version: RedisClient.java adds a parameter to one of its constructors */
 - public RedisClient(final Timer timer, EventLoopGroup group, Class¡? extends SocketChannel¿ socketChannelClass, String host, int port, int connectTimeout, int commandTimeout) {
@@
 As GumTree cannot detect any changes to PackageDeclaration or ImportDeclaration, BuCoR-E cannot extract such edits to suggest related resolutions.
-BuCoR-R is applicable to many cases where (1) BuCoR-E does not work and (2) the needed resolutions are simple or conventional, such as the case in Listing 2.
-However, for unconventional conflicts (see C17–C23 in Table 3) that involve entity deletion, parameter addition/deletion, or type change, BuCoR-R generates nothing because there is no typical, generally accepted way to handle them.
+BuCoR-R is applicable to many cases where (1) BuCoR-E does not work and (2) the needed resolutions are conventional, such as the case in Listing 2.
+However, for conflicts which have no conventional resolution method recommended by literature (see C17–C23 in Table 3), BuCoR-R generates nothing because there is no typical, generally accepted way to handle them.
 For example, Listing 3 shows a conflict from Jedis (jed, 2025).
@@
 4.4. RQ2: BuCoR Resolution Accuracy
-As shown in Table 4, 21 of the 28 resolutions produced by BuCoR-E are correct, leading to 75% accuracy; 29 of the 51 resolutions output by BuCoR-R are correct, achieving 57% accuracy.
-Overall, BuCoR obtained 63% accuracy, having 50 of the 79 generated resolutions to be correct.
-These measurements also show that the two strategies complement each other.
+As shown in Table 4, 21 of the 28 resolutions produced by BuCoR-E are correct, leading to 75% accuracy; 20 of the 51 resolutions output by BuCoR-R are correct, achieving 39% accuracy.
+Overall, BuCoR obtained 52% accuracy, having 41 of the 79 generated resolutions to be correct.
+These measurements also show the two strategies to complement each other.
 Among the 14 cases where both strategies suggest resolutions, BuCoR-E and BuCoR-R correctly resolved 13 and 7 cases, respectively; the correct resolutions overlap in 7 cases.
-Thus, BuCoR resolved 43 cases correctly.
+Thus, BuCoR resolved 34 cases correctly.
 Particularly, BuCoR-E resolved build conflicts by mimicking developers’ project-specific solutions to related build errors.
@@
 Listing 6: A conflict for which the well-accepted resolution pattern does not match developers’ manual resolution (fas, 2025)
-/* Left Version: TypeUtils.java */
+/* Left Version: TypeUtils.java removes an import */
 - import java.beans.Introspector;
-/* Right Version: TypeUtils.java */
+/* Right Version: TypeUtils.java adds a reference to the imported class */
 + if (compatibleWithJavaBean){
 + propertyName= Introspector.decapitalize(methodName.substring(3));
-BuCoR-R sometimes suggested incorrect resolutions, because the frequently applied resolution patterns within it do not always match developers’ practices.
+BuCoR-R suggested incorrect resolutions for three reasons.
+First, its resolution rules do not always match developers’ practices.
 For instance, the conflict in Listing 6 has $l$ remove an import and $r$ add a reference to the removed imported class.
@@
 However, developers implemented a replacement method, to avoid the removed dependency.
-Additionally, for conflicts due to class-hierarchy changes, BuCoR-R modified subclasses to match changes in super classes, while developers sometimes adapted super classes to subclasses.
-The 43 conflicts correctly resolved by BuCoR fall into 15 categories, implying that BuCoR can properly resolve diverse conflicts.
-There are four categories of conflicts for which BuCoR produced some resolutions but none of them are correct: C2, C9, C12, and C18.
-Three of the categories (C2, C9, and C12) are about mismatches between super types and sub types; two of the categories are caused by parameter list changes (C9 and C18).
-Among the seven incorrect resolutions suggested by BuCoR-E, two are partially incorrect, by presenting subsets of the needed changes; five are totally irrelevant.
-Among the 22 incorrect resolutions output by BuCoR-R, 6 are partially incorrect, while 16 are irrelevant.
-Finding 2 (Answer to RQ2): BuCoR generated in total 50 correct resolution(s), achieving 63% (50/79) accuracy. BuCoR-E and BuCoR-R separately contributed 21 and 29 correct resolutions.
+Second, for conflicts due to class-hierarchy changes, BuCoR-R modified subclasses to match changes in super classes, while developers sometimes adapted super classes to subclasses.
+Third, when conflict resolution requires implementation of a new method or updates of a method’s parameter/return type, BuCoR-R can only change the method signature, without customizing method implementation accordingly.
+The 34 conflicts correctly resolved by BuCoR fall into 12 categories, implying that BuCoR can properly resolve diverse conflicts.
+There are seven conflict categories for which BuCoR produced some resolutions but none of them are correct: C2–C4, C8–C9, C12, and C18.
+Six of the categories (C2–C4, C8–C9, and C12) are about mismatches between super types and sub types; three of the categories are caused by parameter list changes (C3, C9 and C18).
+Among the seven incorrect resolutions output by BuCoR-E, two are partially incorrect, by presenting subsets of the needed changes; five are totally irrelevant.
+Among the 31 incorrect resolutions output by BuCoR-R, 15 are partially incorrect, while 16 are irrelevant.
+Finding 2 (Answer to RQ2): BuCoR generated in total 41 correct resolution(s), achieving 52% (41/79) accuracy. BuCoR-E and BuCoR-R separately contributed 21 and 20 correct resolutions.
 4.5. RQ3: BuCoR-E vs. BuCoR-R
@@
 32%) by resolving more conflicts; BuCoR-E achieves higher accuracy by having more of its resolutions match developers’ intents (75% vs.
-57%).
+39%).
 It implies that if developers (1) want to automatically resolve as many conflicts as possible in conventional ways, and (2) have good testing to detect wrong resolutions, they can rely more on BuCoR-R.
@@
 Second, the two strategies are good at resolving different types of conflicts.
-Among the 43 conflicts correctly handled by BuCoR, 7 conflicts are correctly resolved by both strategies.
+Among the 34 conflicts correctly handled by BuCoR, 7 conflicts are correctly resolved by both strategies.
 In addition to that, BuCoR-E correctly resolved 14 cases; for 8 of these cases, BuCoR-R could not suggest anything.
 The eight cases are related to entity removal (i.e., C17, C20, C23) and parameter-list changes (i.e., C21).
-It means that BuCoR-E is better at resolving conflicts in unconventional or complex ways.
-Meanwhile, BuCoR-R correctly resolved 22 cases, for which BuCoR-E suggested nothing.
-These 22 conflicts are mainly related to super-sub mismatches (C3–C4, C8, C10–C11) and duplicated entities (C14, C16).
-It means that BuCoR-R is better at handling simple conflicts in conventional ways.
+It means that BuCoR-E is better at resolving conflicts in unconventional ways.
+Meanwhile, BuCoR-R correctly resolved 13 cases, for which BuCoR-E suggested nothing.
+These 13 conflicts are mainly related to super-sub mismatches (C10–C11) and duplicated entities (C14, C16).
+It means that BuCoR-R is better at handling conflicts in conventional ways.
 4.6. Runtime Overhead
@@
 BuCoR-R includes 16 well-defined rules of handling frequently occurring conflicts.
-Although they cover all the typical conflict-resolution strategies we are aware of, they may not include all the frequently applied strategies in reality.
+Although they cover all the typical conflict-resolution strategies mentioned by prior work (Da Silva et al., 2022; Shen et al., 2022), they may not include all the frequently applied strategies in reality.
 In the future, to enhance our research representativeness, we will (1) expand our dataset to include more merging scenarios, and (2) enlarge the ruleset of BuCoR-R whenever possible.
@@
 However, neither tool pinpoints or resolves build/test conflicts.
-Detectors of Build Conflicts (Wuensche et al., 2020b; Towqir et al., 2023).
+Detectors of Build Conflicts (Wuensche et al., 2020; Towqir et al., 2023).
 The existing tools both statically analyze branch versions, reason about branch edits, and contrast extracted edits with predefined patterns to detect conflicts.
@@
 They are analogous to BuCoR-E, because they all infer and apply patterns by analyzing exemplar edits.
-We did not empirically compare BuCoR with either tool, as they are close-sourced tools to handle conflicts in C/C++ code.
+We did not empirically compare BuCoR with either tool, as they are closed-source and target C/C++ conflicts.
+To assess the effectiveness of large language models (LLMs) on our dataset, we recently conducted an extra experiment.
+We prompted OpenAI’s o4-mini with the full diffs of (base, left) and (base, right) of each conflicting merging scenario.
+Two authors manually validated o4-mini’s outputs.
+Due to token limit, o4-mini was applicable to only 48 scenarios.
+It detected 117 conflicts, of which 63 were correct.
+It correctly resolved 25 of the conflicts—substantially worse than BuCoR.
 BuCoR is more advanced in three aspects.
@@
 6.2. Empirical Studies of Merge Conflicts
-Relationship between Merge Conflicts and Software Maintenance (Estler et al., 2014; Ahmed et al., 2017; Mahmoudi et al., 2019).
+Relationship between Merge Conflicts and Software Maintenance (Estler et al., 2014; Ahmed et al., 2017; Mahmoudi et al., 2019; Vale et al., 2023).
 Estler et al.
@@
 (Mahmoudi et al., 2019) studied the relation between merge conflicts and 15 popular refactoring types.
-These researchers showed that merge conflicts are widespread, although they studied textual conflicts.
+These researchers showed that conflicts are widespread, although they studied textual conflicts.
 Characterization of Merge Conflicts (Ghiotto et al., 2018; Shen et al., 2022; Shen and Meng, 2024; Da Silva et al., 2022).
@@
 In the future, we will improve BuCoR by (i) expanding the rule set of BuCoR-R, and (ii) improving the example mining capability of BuCoR-E when no local example is extractable from branches.
+Acknowledgements.
+We thank all reviewers for their valuable feedback. This work was funded by NSF-1845446.